# Concepts and methods for predicting viral evolution

### v2

@@ 3.1 Tracking viral evolution
Therefore, we track region-specific clade frequencies $x_{r}^{\alpha}(t)$, sublineage frequencies $X_{r}^{\alpha}(t)$, and allele frequencies $x_{r,k,a}(t)$; the frequencies in a given region $r$ are obtained by summation over all strains collected in that region. An inference procedure for empirical fitness trajectories $\hat{f}_{\alpha}(t)$ from region-specific data is discussed in section 3.4. Major surveilled regions are roughly aligned with named WHO transmission zones; region-specific tracking can be refined to individual countries whenever this is relevant for a specific question.
@@ 3.3 Tracking population immunity
may provide a good approximation of the underlying population immunity profile. In the context of predictive analysis, this type of inference has been applied to human antigenic data for SARS-CoV-2 [5]. In this case, individual samples may be assigned a likely recent infection or vaccination dominating cross-immunity; the sample average (15) can then be evaluated as an average over immune classes, equation (13), with weight functions informed by incidence tracking. More generally, the sample average (15) may still give a bona fide input for fitness models; however, detecting and correcting biases in the population sample requires at least partial knowledge of the underlying infection histories by available metadata or computational reconstruction. This practical problem indicates a fundamental challenge of human immune data: given the complexity of human immunization histories [103, 104, 105, 106], only the most common patterns (e.g., the last vaccination or infection) will be represented correctly in realistic population samples. This limits the complexity of immune imprinting and antigenic sin that can be included in data-driven predictive analysis. In the formalism developed here, immune classes define the appropriate coarse-graining of infection and vaccination histories [5]. In practice, we partition human antigenic data into immune classes labelled by one or few previous immunization events, each associated with a specific viral clade, depending on the availability of metadata (Note 3).
@@ 3.4 Evolutionary prediction and validation
the omitted terms are stochastic changes by genetic drift. For sizeable clades of influenza, genetic drift can be neglected, and we obtain a deterministic relation between frequency trajectories and relative fitness that is the central equation of motion used for predictions. Given time-dependent frequency trajectories, we can also compute the change of the mean absolute fitness by evolution, $\Phi(t)=\sum_{\alpha}\dot{x}^{\alpha}(t)F_{\alpha}(t)=\sum_{\alpha}\dot{x}^{\alpha}(t)f_{\alpha}(t)$, the so-called fitness flux [110]. In the deterministic regime, the fitness flux equals the fitness variance in the population of circulating strains, $\Phi(t)={\rm Var}\,F(t)$, which is Fisher’s fundamental theorem.
@@
In a second step, we calibrate the model-based fitness by comparison with empirical fitness data. By equation (4), we obtain clade- and region-specific trajectories of empirical fitness, $\hat{f}_{\alpha,r}(t)$, in a representative set of regions $r$ where clades are in direct competition (see section 3.1). We define a new, calibrated fitness model $F^{c}$ by
@@
The calibration combines clade- and region-specific trajectories of the primary model, here the two-component model $F_{\alpha,r}(t)=F_{\alpha,r}^{\rm ag}(t)+F_{\alpha}^{\rm in}$ computed with region-specific immune weights, and a clade-specific global correction term, $\Delta F_{\alpha}$. Relative fitness computed from the calibrated model, $f^{c}_{\alpha,r}(t)$, is then used as input for predictions.
@@
$S=\sum_{\alpha,r}\int^{t_{0}}(\hat{f}_{\alpha,r}(t)-f^{c}_{\alpha,r}(t))^{2}Y_{r}(t)\,{\rm d}t,$ (23)
@@ 3.5 Vaccine protection
using the nonlinear map (11). This profile applies to a population of otherwise naive hosts, because we single out vaccination as the only source of primary immunization. If the vaccine protection wanes over time (as comprehensively recorded in SARS-CoV-2 [116, 117]), the protection profile $c_{\alpha}^{\rm vac}$ refers to infections shortly after vaccination, after full immune memory has built up but before substantial waning sets in. Given the protection profile for a given vaccine, we can also define the mean protection against circulating strains at a given time,
@@
Importantly, this measure is time-dependent because clade frequencies evolve over time.
@@
$\Delta c^{\rm vac}_{\alpha}(t)=\max\left[c_{\alpha}^{\rm vac}-\bar{C}_{\alpha}(t),0\right],$ (27)
where $c_{\alpha}^{\rm vac}$ is the naive vaccine protection given by equation (25) and $\bar{C}_{\alpha}(t)$ the pre-vaccination population immunity profile given by equation (13) or (15). Here we assume that the immuno-dominant repertoire component activated in response to an infecting strain $\alpha$ is induced by the vaccination or by previous immunizations, whichever component generates the stronger cross-immunity. The resulting differential vaccine protection of individuals, $\Delta c^{{\rm vac},n}_{\alpha}(t)=\max[c_{\alpha}^{\rm vac}-c_{\alpha}^{n},0]$, depends on their immunization history; the population average (27) then follows by equation (15). The underlying immuno-dominance assumption is consistent with a recent biophysical model of lineage activation in an acute infection that predicts a correlation between antigen binding and clone size for antigen-responding antibody lineages [119]. The specific form of complex immune responses remains to be tested in future studies. The resulting differential mean protection against circulating strains,
@@
quantifies the added value of vaccination over population immunity. This measure is time-dependent for two reasons: population immune dynamics affect $\bar{C}_{\alpha}(t)$ and viral evolution propagates the frequencies $x^{\alpha}(t)$. Notably, a good vaccine is expected to decrease in differential protection over time, as population immunity builds up against viral clades $\alpha$ covered by the vaccine.
@@
Importantly, vaccine effectiveness is seen to be a dynamical quantity: it can increase when the naive protection $\bar{c}^{\rm vac}(t)$ increases and decrease when population immunity against clades covered by the vaccine builds up. In a similar way, we can compute estimates of vaccine efficacy as measured in clinical trials, by calibrating the mean protection $\bar{C}(t)$ to known characteristics of the trial cohort.
@@
The vaccine protection measures $\bar{c}^{\rm vac}(t)$ and $\Delta\bar{c}^{\rm vac}(t)$ can be used to rank available candidate vaccines for a given influenza season: higher naive and differential mean protection is predicted to induce lower viral growth in vaccinated population segments without and with previous immunity, respectively. Both measures change over time, underscoring the role of predictive analysis for vaccine strain selection. By using predicted clade frequencies, $x^{\alpha}(t)$, and predicted immune profiles, $\bar{C}_{\alpha}(t)$, we can optimize vaccines in a pre-emptive way, taking into account the likely viral evolution and population immune dynamics up to the season when the vaccine will be used. The protection profiles shown in Fig. 5 and 6 provide a more detailed picture than the mean protection, highlighting viral clades with maximal escape from vaccine-induced and pre-existing population immunity.
@@ 4 Notes
    The established animal model for influenza uses ferrets that are immune-naive against influenza prior to the infection with a reference strain. Neutralization tests of these ferret antisera capture vaccination-induced cross-protection against test strains without confounding effects of pre-existing immunity. The use of ferret data in fitness models for viral evolution in humans depends on a heuristic extrapolation: cross-immunity $c^{\kappa}_{\alpha}$ computed from ferrets infected with a strain from clade $\kappa$ is a good proxi for humans who had their last immunization (infection or vaccination) by a strain from the same clade. An important role of the last immunization in the response to future exposures has been established [103], but older exposures may also be relevant in specific population segments [125, 25]. Another instance of more complex immunisation histories is combinations of a recent infection and vaccination; such cases can be modelled by composite immune classes, $\kappa=(\kappa_{1},\kappa_{2})$ [5]. The impact of immune complexity on viral dynamics has recently been studied in ref. [74].

### v3

@@ 3.1 Tracking viral evolution
Therefore, we track region-specific clade frequencies $x_{r}^{\alpha}(t)$, sublineage frequencies $X_{r}^{\alpha}(t)$, and allele frequencies $x_{r,k,a}(t)$; the frequencies in a given region $r$ are obtained by summation over all strains collected in that region. Major surveilled regions are roughly aligned with named WHO transmission zones; region-specific tracking can be refined to individual countries whenever this is relevant for a specific question. We note that the simple relation (4) between relative frequency change and empirical fitness is valid in a spatially homogeneous population. The regionally diverse seasonal epidemiology of human influenza viruses generates region- and time-dependent absolute growth rates of the viral population. An inference procedure for empirical fitness trajectories $\hat{f}_{\alpha}(t)$ from region-specific data is discussed in section 3.4.
@@ 3.3 Tracking population immunity
may provide a good approximation of the underlying population immunity profile.
@@
Human immune classes
The expressions (13) and (15) for the population immunity profile become equivalent if we partition humans into immune classes representing distinct, suitably coarse-grained infection and vaccination histories [5]. In the context of predictive analysis, class-based inference of immunity profiles has been applied to human antigenic data for SARS-CoV-2 [5]. In this case, individual samples are assigned a likely recent infection or vaccination dominating cross-immunity; the sample average (15) can then be evaluated as an average over immune classes, equation (13), with weight functions informed by incidence tracking. More generally, the sample average (15) may still give a bona fide input for fitness models; however, detecting and correcting biases in the population sample requires at least partial knowledge of the underlying infection histories by available metadata or computational reconstruction. This practical problem indicates a fundamental challenge of human immune data: given the complexity of human immunization histories [104, 105, 106, 107], only the most common patterns (e.g., the last vaccination or infection) will be represented correctly in realistic population samples. This limits the complexity of immune imprinting and antigenic sin that can be included in data-driven predictive analysis. Immune classes address this coarse-graining problem at a tunable level of complexity [5]. In practice, we partition human antigenic data into classes labelled by one or few previous immunization events, each associated with a specific viral clade, depending on the availability of metadata (Note 3).
@@ 3.4 Evolutionary prediction and validation
here we have omitted stochastic changes by genetic drift and correction terms accounting for the spatial heterogeneity of absolute growth rates [5]. For sizeable clades of influenza, genetic drift can be neglected, and we obtain a deterministic relation between frequency trajectories and relative fitness that is the central equation of motion used for predictions. Given time-dependent frequency trajectories, we can also compute the change of the mean absolute fitness by evolution, $\Phi(t)=\sum_{\alpha}\dot{x}^{\alpha}(t)F_{\alpha}(t)=\sum_{\alpha}\dot{x}^{\alpha}(t)f_{\alpha}(t)$, the so-called fitness flux [110]. In the deterministic regime, the fitness flux equals the fitness variance in the population of circulating strains, $\Phi(t)={\rm Var}\,F(t)$, which is Fisher’s fundamental theorem.
@@
In a second step, we calibrate the model-based fitness by comparison with empirical trajectory data $x^{\alpha}_{r}(t)$ in a representative set of regions $r$ where clades are in direct competition (see section 3.1). We define a new, calibrated fitness model $F^{c}$ by
@@
The calibration combines clade- and region-specific trajectories of the primary model, here the two-component model $F_{\alpha,r}(t)=F_{\alpha,r}^{\rm ag}(t)+F_{\alpha}^{\rm in}$ computed with region-specific immune weights, and clade-specific global correction terms $\Delta F_{\alpha}$. The calibrated model inherits the explicit time-dependence of the antigenic fitness component, while the correction terms are taken to be time-independent. Relative fitness computed from the calibrated model, $f^{c}_{\alpha,r}(t)$, is then used as input for predictions.
@@
$S=\int^{t_{0}}\Big{[}\sum_{\alpha,r}(\dot{x}^{\alpha}_{r}(t)-f^{c}_{\alpha,r}(t)x^{\alpha}_{r}(t))Y_{r}(t)\Big{]}^{2}{\rm d}t,$ (23)
@@ 3.5 Vaccine protection
using the nonlinear map (11). This profile applies to a population of otherwise naive hosts, because we single out vaccination as the only source of primary immunization. Given the protection profile of a vaccine, we can also define the mean protection against circulating strains at a given time,
@@
Importantly, this measure is time-dependent because clade frequencies evolve over time. The cross-immunity factors $c_{\alpha}^{\rm vac}$ measure the maximum protection of a given vaccine against strains of a given clade. This value is reached for infections shortly after vaccination, before waning of vaccine-generated humoral immunity sets in. Immune waning generates a time-dependent protection profile $c_{\alpha}^{\rm vac}(t)$, introducing an additional time-dependence in the mean protection $\bar{c}^{\rm vac}(t)$ (Note 7).
@@
$\Delta c^{\rm vac}_{\alpha}(t)=\sum_{\kappa}r_{\kappa}(t)\max[c_{\alpha}^{\rm vac}-c_{\alpha}^{\kappa},0].$ (27)
Here the cross-immunity factors $c_{\alpha}^{\kappa}$ and the immune class weights $r_{k}(t)$ determine pre-existing immunity as in equation (13), and $c_{\alpha}^{\rm vac}$ is the naive vaccine protection given by equation (25). We assume that in a given individual, the immuno-dominant repertoire component activated in response to an infecting strain $\alpha$ is induced by the vaccination or by previous immunizations, whichever component generates the stronger cross-immunity. This immunodominance assumption gives an approximate form for a polyclonal B cell immune response shaped by immune imprinting. It follows from a recent biophysical model of antibody lineage activation in an acute infection that predicts an affinity-graded activation of multiple lineages, generating a correlation between antigen affinity and clone size [117]. The resulting differential protection profile is a host population average of immunodominance, $\Delta c^{\rm vac}_{\alpha}(t)=\frac{1}{N}\sum_{n=1}^{N}\max[c_{\alpha}^{\rm vac}-c_{\alpha}^{n},0]$; the form (27) then follows by grouping individuals into immune classes. We can also evaluate the differential mean protection against circulating strains,
@@
which quantifies the added value of vaccination over population immunity. This measure is time-dependent for two reasons: population immune dynamics affect $\bar{C}_{\alpha}(t)$ and viral evolution propagates the frequencies $x^{\alpha}(t)$. Notably, a good vaccine is expected to decrease in differential protection over time, as population immunity builds up against viral clades $\alpha$ covered by the vaccine.
@@
Importantly, vaccine effectiveness is seen to be a dynamical quantity: it can increase when the naive protection $\bar{c}^{\rm vac}(t)$ increases and decrease when population immunity against clades covered by the vaccine builds up. In a similar way, we can compute vaccine efficacy as measured in clinical trials, by calibrating the mean protection $\bar{C}(t)$ to known characteristics of the trial cohort. To include the waning of vaccine-induced immunity, we evaluate equations (27) and (28) with a time-dependent vaccine protection profile $c_{\alpha}^{\rm vac}(t)$ (Note 7). These predictive VE estimates can be computed from molecular and surveillance data available before a given influenza season. They do not take into account external factors such as timing and severity the seasonal epidemic. Relative vaccine effectiveness, here differences between candidate vaccines based on different vaccine strains, is expected to be less sensitive to these external factors than the absolute effectiveness of a given vaccine. Therefore, the main use of predictive VE estimates is the pre-emptive ranking of candidate vaccines, as described in the next subsection.
@@
The vaccine protection measures $\bar{c}^{\rm vac}(t)$ and $\Delta\bar{c}^{\rm vac}(t)$ can be used to rank available candidate vaccines for a given influenza season. By equation (29), the ranking by $\Delta\bar{c}^{\rm vac}(t)$ is equivalent to a ranking by VE. Importantly, vaccine protection changes over time due to viral evolution and population immunity dynamics, underscoring the role of predictive analysis for vaccine strain selection. By using predicted clade frequencies, $x^{\alpha}(t)$, and predicted immune profiles, $\bar{C}_{\alpha}(t)$, we can optimize vaccines in a pre-emptive way, taking into account the likely viral evolution and population immune dynamics up to and throughout the season when the vaccine will be used. In addition, we can account for the partial waning of vaccine protection between vaccination and the expected next exposure (Note 7). The protection profiles shown in Fig. 5 and 6 provide a more detailed picture than the mean protection, highlighting viral clades with maximal escape from vaccine-induced and pre-existing population immunity.
@@ 4 Notes
    The established animal model for influenza uses ferrets that are immune-naive against influenza prior to the infection with a reference strain. Neutralization tests of these ferret antisera capture vaccination-induced cross-protection against test strains without confounding effects of pre-existing immunity. The use of ferret data in fitness models for viral evolution in humans depends on a heuristic extrapolation: cross-immunity $c^{\kappa}_{\alpha}$ computed from ferrets infected with a strain from clade $\kappa$ is a good proxi for humans who had their last immunization (infection or vaccination) by a strain from the same clade. More complex immunization histories are, for example, combinations of a recent infection and vaccination; such cases can be modelled by composite immune classes, $\kappa=(\kappa_{1},\kappa_{2})$ [5]. An important role of the last immunization in the response to future exposures has been established [104]. Older exposures may also be relevant in specific population segments [123, 25]. The impact of immune complexity on viral dynamics has recently been studied in ref. [74].